# Multi-View Adaptive Contrastive Learning for Information Retrieval Based Fault Localization

### v2

@@ 5.1.1 Dataset
To evaluate the effectiveness of our proposed approach, we use the same benchmark datasets provided by Ye et al. [6], which are widely used in fault localization research works [5, 7, 9, 10, 11, 12, 14, 13, 15, 29, 16]. The datasets have five real-world projects (Birt, Eclipse Platform UI, JDT, SWT, Tomcat), which contain ground truth associated with each bug report. For comparison, we follow previous studies to evaluate our approach to each project separately.
Table 2 shows the statistics of the datasets. The column of #Bug Reports shows the number of bug reports of each project. In line with Qi et al. [13], since some buggy files may be modified or deleted in the fixed version, for each bug we check out the source code files of the prefix version, and remove the invalid buggy paths which do not exist in the prefix version. After that, we filter out bug reports that without linking any buggy paths. Furthermore, we filter out bug reports from the validation and test sets whose ground truth (i.e., relevant buggy files) are not present in the training set. The values in parentheses are the number of bug reports we filtered. The column of #Source Files shows the number of source code file type nodes in report-code graph and code-code graph. Specifically, we chronologically sort bug reports by commit timestamp and choose the oldest version as the first version. We keep all the source code files from the first version as nodes in the network. Since the number of source code files in the next version may be changed by adding or deleting files, we perform the following steps: 1) for newly added files, we directly add them to the network; 2) for deleted files, we do not remove them from the network, since they may be linked to bug reports from the previous version; 3) for modified files, if only the contents of the file are modified, the nodes in the network do not make any changes. If the name of the modified file is different from the previous version, we add the modified file as a new node to the network. Follow these steps until the latest version in training set has been traversed, and the final number of source code files of each project is shown in column #Source Files. The column of #Buggy Files shows the number of fixed buggy files, and #Buggy Rate indicates the percentage of buggy files within each project.
@@
Table 2: Statistics of benchmark dataset. ##\##Bug Reports represents the number of bug reports of each project, and the values in parentheses represent the actual number of bug reports filtered in this paper.
Project Time Range ##\##Bug Reports ##\##Source Files ##\##Buggy Files ##\##Buggy Rate
@@
Table 3: Statistics of three types of interactions of each project. ##\##RC, ##\##RR and ##\##CC denote the number of report-code relationships, report-report relationships and code-code relationships, respectively.
Project ##\##RC ##\##RR ##\##CC Avg.##\##RC per Avg.##\##RR per Avg.##\##CC per
@@ 5.1.2 Evaluation metrics
To evaluate our approach, we use Accuracy@N, mean average precision (MAP) [56], and mean reciprocal rank (MRR) [57] as the evaluation metrics, which are commonly used in software engineering. The higher the Accuracy@N, MRR, and MAP values, the better the fault localization performance.
Accuracy@N is the percentage of bug reports whose at least one buggy file is ranked in the top $N$ files in the recommended list. It measures the overall ranking performance of the localization model, which is broadly used in prior studies [13]. We consider three values of $N$: 1, 5, and 10.
@@
Moreover, we use the Wilcoxon signed rank test [58] to check whether the performance difference between our model with baselines is significant. We consider that one approach performs significantly better than the other one at the 95% confidence level if the corresponding $p$-value is less than 0.05, otherwise, there is no statistically significant difference between the two approaches. We also employ Cliff’s delta ($\delta$) [59] to quantify the amount of difference between two approaches. Here, the effectiveness level is considered negligible if $|\delta|<0.147$, small if $0.330.147\leq|\delta|<$ 0.33, medium if $0.4740.33\leq|\delta|<$ 0.474, and large if $|\delta|\geq 0.474$.
@@ 5.1.3 Baselines
1) BugLocator [5]: an IR-based approach that utilized revised vector space model (rVSM) to calculate the text similarity between bug reports and source code files, and then ranked all the source code files to locate potential buggy files to a given bug report.
2) DNNLOC [9]: a deep learning approach that combined the features built from deep neural network (DNN), rVSM, and project’s bug-fixing history.
3) KGBugLocator [14]: a knowledge graph and deep learning approach that constructed code knowledge graph to enrich semantic information of code for localization.
4) DreamLoc [13]: a deep learning approach that employed the attention mechanism to calculate the relevance between bug report terms and code snippets.
5) FLIM [15]: a deep learning approach that utilized a large language model (i.e., CodeBERT) to extract semantic features of source code file at the function-level.
6) BL-GAN [29]: a semi-supervised learning approach that employed generative adversarial network to capture the potential relevance distribution between bug report and source code file from both limited bug-fix records and abundant unfixed bug reports.
7) bjXnet [16]: a deep learning approach that combined the text semantic features of the source code and the graph features of the code property graph to enrich semantic information of code for localization.
@@ 5.1.4 Implementation details
We use pytorch to construct the model and all experiments are run on Windows Server 2016 Datacenter with 6 cores of 2.3 GHz CPU, 48 GB RAM and NVIDIA Tesla V100 GPUs with 32 GB memory. We use the uniform distributed random initializer to generate the initial embeddings of bug reports and source code files. GNN training is at full batch, and the embedding size defaults to 32 for each node in the three views. Here, we optimize our model with the Adam optimizer [60] with learning rate 0.01 for training. For regularization, we apply the early stopping strategy [61] for training. In all experiments, the optimal number of GNN layers is searched in $\{1,2,3,4,5\}$, which is discussed in detail in RQ3 of Section 5.2. For the projects Birt, JDT, and Tomcat, the number of layers is set to 4, while for Eclipse Platform UI and SWT, is set to 3. The temperature of contrastive learning $\tau$ is set to 0.1. To avoid over-fitting, we dropout (dropout rate is 0.2) in every layer of our proposed MACL-IRFL. We split each dataset into a training dataset containing 80% data, a validation dataset containing 10% and a test dataset containing 10% data by chronological order.
@@ 5.3 Experimental Results for Research Questions
Table 6: Statistical test results of our approach vs. other approaches based on Wilcoxon signed-rank test and Cliff delta, where avg denotes the average result of each approach on five projects. L and M denote the large (i.e., $|\delta|\geq 0.474$) and medium (i.e., $0.4740.33\leq|\delta|<$ 0.474) difference between our model and other approaches.
@@
Figure 8: Two dimensional projections of t-SNE [62] visualization embeddings of different variants for Tomcat dataset. The red nodes and bule nodes denote the distribution of source code files and bug reports respectively.
@@
First, when the layers $L$ increase from 1 to 3, the performance of MACL-IRFL can be improved substantially. In particular, MACL-IRFL-2 and MACL-IRFL-3 achieve consistent an improvement over MACL-IRFL-1 on all datasets. For example, MACL-IRFL-2 outperforms MACL-IRFL-1 by 23.81%, 28.37%, 30.33%, 22.89% and 25.71% on Accuracy@1, Accuracy@5, Accuracy@10, MAP, and MRR on the five datasets, respectively. MACL-IRFL-3 outperforms MACL-IRFL-1 by 39.94%, 48.87%, 48.99%, 39.32% and 43.16% on Accuracy@1, Accuracy@5, Accuracy@10, MAP, and MRR on the five datasets, respectively. Moreover, the statistical test results also show that both MACL-IRFL-2 and MACL-IRFL-3 perform significantly better than MACL-IRFL-1 ($p$-value = 0.0079 <0.05, and $|\delta|$ = 1.0 $\geq$ 0.474). These improvements mainly benefit from the modeling of the high-order relationships among bug reports and source code files, carried by the high-layer GNN embedding propagation. This validates the conclusion of previous studies [36] that nodes incrementally gain more and more information from further reaches of the graph as aggregation layers or search depth deepens.
@@ 6 Threats to Validity
One of the key threats to internal validity relates to author bias in the experiments. The dataset that we obtain is taken from prior papers [11, 12], and has been evaluated by other fault localization techniques [5, 7, 9, 10, 11, 12, 14, 13, 15, 29, 16] The data are bug reports taken from bug tracking system from real projects and thus are realistic. Another key threat is the specific parameter choices we used to build our MACL-IRFL model. To mitigate the impact of this threat, all parameters were either studied by us or were reported in other prior reputable papers as recommended or optimal [5, 13, 29]. Therefore, we believe there are limited threats to the internal validity of this study.
@@
The threats to construct validity relate to the suitability of our evaluation metrics. We have used Accuracy@N, MAP, and MRR as evaluation metrics. These metrics were also used by prior fault localization studies [5, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], thereby mitigating concerns regarding their suitability and suggesting limited threats to construct validity. Furthermore, MACL-IRFL is based on GNN. One of the limitations of GNN-based approaches is the scalability [63]. Specifically, when dealing with large graphs with a high number of nodes and intricate relationships, GNN-based approaches will encounter challenges due to substantial memory requirements and protracted training durations. To mitigate this threat, we plan to explore subsampling strategies to make our model more applicable in future work.
@@ 7.1 Text-based fault localization
The topic-based model assumes that both bug reports and source code are generated by some topic, thereby enabling the similarity between bug reports and source code to be calculated based on their topic distributions. Lukins et al. [20, 21] applied Latent Dirichlet Allocation (LDA) to extract latent topics from source code files and bug reports, subsequently indexing source files based on these topics. Given a new bug report, they constructed a textual query from its description and searched the indexed source files using the VSM. Nguyen et al. [22] introduced BugScout, which used a specialized LDA-based topic model to represent technical aspects as topics in bug reports and source files. BugScout comprises two components: the S-component models source files, while the B-component, an extended LDA model, considers both a bug report’s own topic distribution and the topic distributions of its corresponding buggy source files. It predicts the relevant source files for a new bug report by estimating its topic distribution and comparing it with historical source file distributions, prioritizing files with a history of more defects.
@@
The Vector Space Model (VSM) is the most frequently utilized by many fault localization techniques. VSM is based on the bag-of-words model, which regards bug reports and souce code files as document. Each document is expressed as a vector of token weights typically computed as a product of TF (Term frequency) and IDF (Inverse Document Frequency) of each token. Subsequently, the cosine similarity is used to determine how close the two vectors are. For instance, Zhou et al. [5] proposed BugLocator, which utilized rVSM and considered information about similar bugs that have been fixed before. Firstly, they employed rVSM to search and rank relevant source code files. Subsequently, they collected the similar bugs that have been fixed before, and ranked the relevant files by analyzing past similar bugs and their fixes. Ultimately, they combined the ranks from both sources, utilizing a final score to present a prioritized list of source files, thereby facilitating the efficient localization of new bugs. Saha et al. [23] proposed the BLUiR approach, which divides bug reports into two parts: summary and description, and segments source code into four components: class names, method names, variable names, and comments. Based on a pre-trained BM25 model, these segments are represented as feature vectors. The textual similarity scores between the bug report and each component of the source code are then calculated individually and aggregated to form the final textual similarity score.
@@
Some researchers transformed the fault localization task into a ranking problem based on parameter learning from an optimization function perspective, specifically optimizing the ranking of the relevance between current source code and a given bug report according to the historical bug report resolution information. Ye et al. [6] introduced an adaptive ranking approach using a learning-to-rank technique and leveraged domain knowledge features in software engineering for locating relevant buggy files, such as API descriptions of library components used in the code, the bug-fixing history, and the code change history. They extracted six features (surface lexical similarity, API-enriched lexical similarity, collaborative filtering score, bug-fixing recency, and bug-fixing frequency) and gave scores to source files by the weighted combination of these features. Gharibi et al. [10] presented a multi-component fault localization approach that leveraged different textual properties of bug reports and source files as well as relations between previously fixed bug reports and a newly received one. Their approach has five components that are the token matching, VSM similarity, stack trace, semantic similarity, and fixed bug reports. Each of the mentioned components gives a score to source files with respect to each bug report. Then these scores are combined to get a final ranking score that is used to rank relevant source files.
@@ 7.2 Semantic-based fault localization
Huo et al. [64] proposed NP-CNN, which leveraged Convolutional Neural Networks (CNNs) to learn unified features from natural language and source code in programming language for bug localization problems. Their approach includes intra-language feature extraction layers and cross-language feature fusion layers. The former employed separate CNNs to extract intra-language features form natural language and programming language, while the latter employed a fully connected neural network to fuse middle-level features extracted from bug reports and source code files to generate a unified feature representation. Finally, the unified feature representation is then passed through a fully connected layer for classification.
@@
Lam et al. [8, 9], proposed DNNLOC, which combines a kind of deep neural network (Auto-Encoder) and the rVSM to improve bug localization performance. Their approach firstly employed rVSM to capture textual similarity features between bug reports and source code. It further learned the associations between identifiers, API method calls and classes, comments, and API descriptions in the code within bug reports and source code. Finally, a feature fusion layer is utilized to learn the weights to combine all features for an accurate ranking score.
@@
Ye et al. [7] first leveraged word embedding to project text and code as meaning vectors into a shared representation space to bridge the lexical gap between natural language text and source code. In the proposed architecture, word embeddings are first trained on API documents, tutorials, and reference documents, and then aggregated in order to estimate semantic similarities between documents.
@@
Xiao et al. [11] proposed a deep learning model named DeepLocator, which consisted of an enhanced CNN together with a new rTF-IDuF method and the pre-trained word2vec technique to improve the performance of fault localization. The follow-up work DeepLoc of Xiao et al. [12] used two different word-embedding techniques (Sent2Vec and combined word2vec) to convert bug reports and source files into vectors respectively. Xiao et al. [65] further employed CNN with multiple filters and an ensemble of random forests with multi-grained scanning to extract semantic and structural features from the word vectors derived from bug reports and source files. Finally, a cascade forest consisting of two Completely Random Tree Forests (CRTFs) and two random forests is employed to extract deep features from both bug reports and source code files and perform classification.
@@
Qi et al. [13] proposed a framework DreamLoc, which utilized a relevance matching model to locate buggy files. DreamLoc comprises of three components, the Wide component measured the relevance from the perspective of domain-specific features in software engineering, while the Deep component is a deep relevance matching model, which combined local matching and global matching and considers the exact matching of keywords, the local details in source files as well as the difference in importance of words. Finally, the Fusion component fused the results of Wide and Deep components and gets the final output that shows how likely the source file is relevant to the bug report.
@@
Huo et al. [28] proposed CG-CNN, which enhanced the unified features for bug localization by exploiting structural and sequential nature from the control flow graph. Specifically, CG-CNN firstly used CNN and DeepWalk model to learn feature representation of each statement and further extract semantics from control flow graph based on multi-instance decomposition.
@@
Zhang et al. [14] proposed KGBugLocator, which also extracted rich information of source code from the perspective of graph. KGBugLocator solved the fault localization problem via code knowledge graph embeddings, code-specific features and bug-specific features. Specifically, KGBugLocator first exploited interrelation information via code knowledge graph. Subsequently, it employed a keywords supervised bi-directional attention mechanism to mine the interaction information between bug reports and source code files. Ultimately, a fault localization component calculated similarity scores based on the integrated representations.
@@
Liang et al. [15] proposed FLIM, which leveraged CodeBERT to capture function-level code semantics and computed bug report-source code similarity through aggregated function interactions. FLIM consists of two stages. In the first stage, they initially built a training dataset called MIX dataset by splitting a source code file into a set of functions. Following this, they utilized the MIX dataset to fine-tune the CodeBERT model to make it adaptive to fault localization task. In the second stage, they employed the fine-tuned CodeBERT model to predict the relevance between the functions of the source code files and two parts (i.e., summary and description) of the bug report. Finally, semantic features extracted from CodeBERT and IR features are fused to a learning-to-rank model which gives the final relevance score.
@@
Han et al. [16] proposed a multi-modal representation learning framework bjXnet, which integrated the text semantic features of the source code (shallow semantic features) and the graph features of the CPG (deep semantic features) by using the attention mechanism. Specifically, bjXnet employed the GatedGraphConv layer based on the GGNN to build the GatedGCN network for graph feature extraction, and utilized TextCNN as the text feature extractor to extract the text content of bug reports and source code files. Finally, bjXnet embedded bug reports, source code text, and source code CPGs into the same semantic space, comparing the relationship between report representation vectors and code representation vectors by the cosine similarity to determine whether the report and source code are related.
@@
Apart from focusing solely on semantic matching techniques, some researchers have embarked on endeavors to improve the performance of fault localization by exploring other perspectives and methods combined with advanced semantic models. For instances, Zhu et al. [29] proposed BL-GAN, which leveraged the adversarial training strategy and reinforcement learning for fault localization. In specific, the generator in their model generated code file paths for each given bug report, so that it can further construct pairs of bug report and code file. On the other hand, the discriminator attempted to distinguish whether the pair of bug report and code file is from the generator or not. BL-GAN adopted an attention-based Transformer architecture to capture semantic and sequence information to process bug reports, while incorporated a novel multi-layer GCN to process the source code in a graphical view to capture the proprietary structural information in code files, thereby enhancing the performance of fault localization.

### v3

@@ 5.1.1 Dataset
To evaluate the effectiveness of our proposed approach, we use the same benchmark datasets provided by Ye et al. [ye2014learning], which are widely used in fault localization research works [zhou2012should, ye2016word, lam2017bug, gharibi2018leveraging, xiao2017improving, xiao2019improving, zhang2020exploiting, qi2021dreamloc, liang2022modeling, zhu2022bl, han2023bjxnet]. The datasets have five real-world projects (Birt, Eclipse Platform UI, JDT, SWT, Tomcat), which contain ground truth associated with each bug report. For comparison, we follow previous studies to evaluate our approach to each project separately.
Table 2 shows the statistics of the datasets. The column of #Bug Reports shows the number of bug reports of each project. In line with Qi et al. [qi2021dreamloc], since some buggy files may be modified or deleted in the fixed version, for each bug we check out the source code files of the prefix version, and remove the invalid buggy paths which do not exist in the prefix version. After that, we filter out bug reports that without linking any buggy paths. Furthermore, we filter out bug reports from the validation and test sets whose ground truth (i.e., relevant buggy files) are not present in the training set. The values in parentheses are the number of bug reports we filtered. The column of #Source Files shows the number of source code file type nodes in report-code graph and code-code graph. Specifically, we chronologically sort bug reports by commit timestamp and choose the oldest version as the first version. We keep all the source code files from the first version as nodes in the network. Since the number of source code files in the next version may be changed by adding or deleting files, we perform the following steps: 1) for newly added files, we directly add them to the network; 2) for deleted files, we do not remove them from the network, since they may be linked to bug reports from the previous version; 3) for modified files, if only the contents of the file are modified, the nodes in the network do not make any changes. If the name of the modified file is different from the previous version, we add the modified file as a new node to the network. Follow these steps until the latest version in training set has been traversed, and the final number of source code files of each project is shown in column #Source Files. The column of #Buggy Files shows the number of fixed buggy files, and #Buggy Rate indicates the percentage of buggy files within each project.
@@
Table 2: Statistics of benchmark dataset. #\#Bug Reports represents the number of bug reports of each project, and the values in parentheses represent the actual number of bug reports filtered in this paper.
Project Time Range #\#Bug Reports #\#Source Files #\#Buggy Files #\#Buggy Rate
@@
Table 3: Statistics of three types of interactions of each project. #\#RC, #\#$R$ and #\#$C$ denote the number of report-code relationships, report-report relationships and code-code relationships, respectively.
Project #\#RC #\#$R$ #\#$C$ Avg.#\#RC per Avg.#\#$R$ per Avg.#\#$C$ per
@@ 5.1.2 Evaluation metrics
To evaluate our approach, we use Accuracy@N, mean average precision (MAP) [schutze2008introduction], and mean reciprocal rank (MRR) [voorhees2001trec] as the evaluation metrics, which are commonly used in software engineering. The higher the Accuracy@N, MRR, and MAP values, the better the fault localization performance.
Accuracy@N is the percentage of bug reports whose at least one buggy file is ranked in the top $N$ files in the recommended list. It measures the overall ranking performance of the localization model, which is broadly used in prior studies [qi2021dreamloc]. We consider three values of $N$: 1, 5, and 10.
@@
Moreover, we use the Wilcoxon signed rank test [wilcoxon1970critical] to check whether the performance difference between our model with baselines is significant. We consider that one approach performs significantly better than the other one at the 95% confidence level if the corresponding $p$-value is less than 0.05, otherwise, there is no statistically significant difference between the two approaches. We also employ Cliff’s delta ($\delta$) [macbeth2011cliff] to quantify the amount of difference between two approaches. Here, the effectiveness level is considered negligible if $|\delta|<0.147$, small if $0.147\leq|\delta|<0.33$, medium if $0.33\leq|\delta|<0.474$, and large if $|\delta|\geq 0.474$.
@@ 5.1.3 Baselines
1) BugLocator [zhou2012should]: an IR-based approach that utilized revised vector space model (rVSM) to calculate the text similarity between bug reports and source code files, and then ranked all the source code files to locate potential buggy files to a given bug report.
2) DNNLOC [lam2017bug]: a deep learning approach that combined the features built from deep neural network (DNN), rVSM, and project’s bug-fixing history.
3) KGBugLocator [zhang2020exploiting]: a knowledge graph and deep learning approach that constructed code knowledge graph to enrich semantic information of code for localization.
4) DreamLoc [qi2021dreamloc]: a deep learning approach that employed the attention mechanism to calculate the relevance between bug report terms and code snippets.
5) FLIM [liang2022modeling]: a deep learning approach that utilized a large language model (i.e., CodeBERT) to extract semantic features of source code file at the function-level.
6) BL-GAN [zhu2022bl]: a semi-supervised learning approach that employed generative adversarial network to capture the potential relevance distribution between bug report and source code file from both limited bug-fix records and abundant unfixed bug reports.
7) bjXnet [han2023bjxnet]: a deep learning approach that combined the text semantic features of the source code and the graph features of the code property graph to enrich semantic information of code for localization.
@@ 5.1.4 Implementation details
We use pytorch to construct the model and all experiments are run on Windows Server 2016 Datacenter with 6 cores of 2.3 GHz CPU, 48 GB RAM and NVIDIA Tesla V100 GPUs with 32 GB memory. We use the uniform distributed random initializer to generate the initial embeddings of bug reports and source code files. GNN training is at full batch, and the embedding size defaults to 32 for each node in the three views. Here, we optimize our model with the Adam optimizer [kingma2014adam] with learning rate 0.01 for training. For regularization, we apply the early stopping strategy [caruana2000overfitting] for training. In all experiments, the optimal number of GNN layers is searched in $\{1,2,3,4,5\}$, which is discussed in detail in RQ3 of Section 5.2. For the projects Birt, JDT, and Tomcat, the number of layers is set to 4, while for Eclipse Platform UI and SWT, is set to 3. The temperature of contrastive learning $\tau$ is set to 0.1. To avoid over-fitting, we dropout (dropout rate is 0.2) in every layer of our proposed MACL-IRFL. We split each dataset into a training dataset containing 80% data, a validation dataset containing 10% and a test dataset containing 10% data by chronological order.
@@ 5.3 Experimental Results for Research Questions
Table 6: Statistical test results of our approach vs. other approaches based on Wilcoxon signed-rank test and Cliff delta, where avg denotes the average result of each approach on five projects. L and M denote the large (i.e., $|\delta|\geq 0.474$) and medium (i.e., $0.33\leq|\delta|<0.474$) difference between our model and other approaches.
@@
Figure 8: Two dimensional projections of t-SNE [van2008visualizing] visualization embeddings of different variants for Tomcat dataset. The red nodes and bule nodes denote the distribution of source code files and bug reports respectively.
@@
First, when the layers $L$ increase from 1 to 3, the performance of MACL-IRFL can be improved substantially. In particular, MACL-IRFL-2 and MACL-IRFL-3 achieve consistent an improvement over MACL-IRFL-1 on all datasets. For example, MACL-IRFL-2 outperforms MACL-IRFL-1 by 23.81%, 28.37%, 30.33%, 22.89% and 25.71% on Accuracy@1, Accuracy@5, Accuracy@10, MAP, and MRR on the five datasets, respectively. MACL-IRFL-3 outperforms MACL-IRFL-1 by 39.94%, 48.87%, 48.99%, 39.32% and 43.16% on Accuracy@1, Accuracy@5, Accuracy@10, MAP, and MRR on the five datasets, respectively. Moreover, the statistical test results also show that both MACL-IRFL-2 and MACL-IRFL-3 perform significantly better than MACL-IRFL-1 ($p$-value = 0.0079 <0.05, and $|\delta|$ = 1.0 $\geq$ 0.474). These improvements mainly benefit from the modeling of the high-order relationships among bug reports and source code files, carried by the high-layer GNN embedding propagation. This validates the conclusion of previous studies [hamilton2017inductive] that nodes incrementally gain more and more information from further reaches of the graph as aggregation layers or search depth deepens.
@@ 6 Threats to Validity
One of the key threats to internal validity relates to author bias in the experiments. The dataset that we obtain is taken from prior papers [xiao2017improving, xiao2019improving], and has been evaluated by other fault localization techniques [zhou2012should, ye2016word, lam2017bug, gharibi2018leveraging, xiao2017improving, xiao2019improving, zhang2020exploiting, qi2021dreamloc, liang2022modeling, zhu2022bl, han2023bjxnet] The data are bug reports taken from bug tracking system from real projects and thus are realistic. Another key threat is the specific parameter choices we used to build our MACL-IRFL model. To mitigate the impact of this threat, all parameters were either studied by us or were reported in other prior reputable papers as recommended or optimal [zhou2012should, qi2021dreamloc, zhu2022bl]. Therefore, we believe there are limited threats to the internal validity of this study.
@@
The threats to construct validity relate to the suitability of our evaluation metrics. We have used Accuracy@N, MAP, and MRR as evaluation metrics. These metrics were also used by prior fault localization studies [zhou2012should, ye2016word, lam2015combining, lam2017bug, gharibi2018leveraging, xiao2017improving, xiao2019improving, qi2021dreamloc, zhang2020exploiting, liang2022modeling, han2023bjxnet], thereby mitigating concerns regarding their suitability and suggesting limited threats to construct validity. Furthermore, MACL-IRFL is based on GNN. One of the limitations of GNN-based approaches is the scalability [fan2017big]. Specifically, when dealing with large graphs with a high number of nodes and intricate relationships, GNN-based approaches will encounter challenges due to substantial memory requirements and protracted training durations. To mitigate this threat, we plan to explore subsampling strategies to make our model more applicable in future work.
@@ 7.1 Text-based fault localization
The topic-based model assumes that both bug reports and source code are generated by some topic, thereby enabling the similarity between bug reports and source code to be calculated based on their topic distributions. Lukins et al. [lukins2008source, lukins2010bug] applied Latent Dirichlet Allocation (LDA) to extract latent topics from source code files and bug reports, subsequently indexing source files based on these topics. Given a new bug report, they constructed a textual query from its description and searched the indexed source files using the VSM. Nguyen et al. [nguyen2011topic] introduced BugScout, which used a specialized LDA-based topic model to represent technical aspects as topics in bug reports and source files. BugScout comprises two components: the S-component models source files, while the B-component, an extended LDA model, considers both a bug report’s own topic distribution and the topic distributions of its corresponding buggy source files. It predicts the relevant source files for a new bug report by estimating its topic distribution and comparing it with historical source file distributions, prioritizing files with a history of more defects.
@@
The Vector Space Model (VSM) is the most frequently utilized by many fault localization techniques. VSM is based on the bag-of-words model, which regards bug reports and souce code files as document. Each document is expressed as a vector of token weights typically computed as a product of TF (Term frequency) and IDF (Inverse Document Frequency) of each token. Subsequently, the cosine similarity is used to determine how close the two vectors are. For instance, Zhou et al. [zhou2012should] proposed BugLocator, which utilized rVSM and considered information about similar bugs that have been fixed before. Firstly, they employed rVSM to search and rank relevant source code files. Subsequently, they collected the similar bugs that have been fixed before, and ranked the relevant files by analyzing past similar bugs and their fixes. Ultimately, they combined the ranks from both sources, utilizing a final score to present a prioritized list of source files, thereby facilitating the efficient localization of new bugs. Saha et al. [saha2013improving] proposed the BLUiR approach, which divides bug reports into two parts: summary and description, and segments source code into four components: class names, method names, variable names, and comments. Based on a pre-trained BM25 model, these segments are represented as feature vectors. The textual similarity scores between the bug report and each component of the source code are then calculated individually and aggregated to form the final textual similarity score.
@@
Some researchers transformed the fault localization task into a ranking problem based on parameter learning from an optimization function perspective, specifically optimizing the ranking of the relevance between current source code and a given bug report according to the historical bug report resolution information. Ye et al. [ye2014learning] introduced an adaptive ranking approach using a learning-to-rank technique and leveraged domain knowledge features in software engineering for locating relevant buggy files, such as API descriptions of library components used in the code, the bug-fixing history, and the code change history. They extracted six features (surface lexical similarity, API-enriched lexical similarity, collaborative filtering score, bug-fixing recency, and bug-fixing frequency) and gave scores to source files by the weighted combination of these features. Gharibi et al. [gharibi2018leveraging] presented a multi-component fault localization approach that leveraged different textual properties of bug reports and source files as well as relations between previously fixed bug reports and a newly received one. Their approach has five components that are the token matching, VSM similarity, stack trace, semantic similarity, and fixed bug reports. Each of the mentioned components gives a score to source files with respect to each bug report. Then these scores are combined to get a final ranking score that is used to rank relevant source files.
@@ 7.2 Semantic-based fault localization
Huo et al. [huo2016learning] proposed NP-CNN, which leveraged Convolutional Neural Networks (CNNs) to learn unified features from natural language and source code in programming language for bug localization problems. Their approach includes intra-language feature extraction layers and cross-language feature fusion layers. The former employed separate CNNs to extract intra-language features form natural language and programming language, while the latter employed a fully connected neural network to fuse middle-level features extracted from bug reports and source code files to generate a unified feature representation. Finally, the unified feature representation is then passed through a fully connected layer for classification.
@@
Lam et al. [lam2015combining, lam2017bug], proposed DNNLOC, which combines a kind of deep neural network (Auto-Encoder) and the rVSM to improve bug localization performance. Their approach firstly employed rVSM to capture textual similarity features between bug reports and source code. It further learned the associations between identifiers, API method calls and classes, comments, and API descriptions in the code within bug reports and source code. Finally, a feature fusion layer is utilized to learn the weights to combine all features for an accurate ranking score.
@@
Ye et al. [ye2016word] first leveraged word embedding to project text and code as meaning vectors into a shared representation space to bridge the lexical gap between natural language text and source code. In the proposed architecture, word embeddings are first trained on API documents, tutorials, and reference documents, and then aggregated in order to estimate semantic similarities between documents.
@@
Xiao et al. [xiao2017improving] proposed a deep learning model named DeepLocator, which consisted of an enhanced CNN together with a new rTF-IDuF method and the pre-trained word2vec technique to improve the performance of fault localization. The follow-up work DeepLoc of Xiao et al. [xiao2019improving] used two different word-embedding techniques (Sent2Vec and combined word2vec) to convert bug reports and source files into vectors respectively. Xiao et al. [xiao2018bug] further employed CNN with multiple filters and an ensemble of random forests with multi-grained scanning to extract semantic and structural features from the word vectors derived from bug reports and source files. Finally, a cascade forest consisting of two Completely Random Tree Forests (CRTFs) and two random forests is employed to extract deep features from both bug reports and source code files and perform classification.
@@
Qi et al. [qi2021dreamloc] proposed a framework DreamLoc, which utilized a relevance matching model to locate buggy files. DreamLoc comprises of three components, the Wide component measured the relevance from the perspective of domain-specific features in software engineering, while the Deep component is a deep relevance matching model, which combined local matching and global matching and considers the exact matching of keywords, the local details in source files as well as the difference in importance of words. Finally, the Fusion component fused the results of Wide and Deep components and gets the final output that shows how likely the source file is relevant to the bug report.
@@
Huo et al. [huo2020control] proposed CG-CNN, which enhanced the unified features for bug localization by exploiting structural and sequential nature from the control flow graph. Specifically, CG-CNN firstly used CNN and DeepWalk model to learn feature representation of each statement and further extract semantics from control flow graph based on multi-instance decomposition.
@@
Zhang et al. [zhang2020exploiting] proposed KGBugLocator, which also extracted rich information of source code from the perspective of graph. KGBugLocator solved the fault localization problem via code knowledge graph embeddings, code-specific features and bug-specific features. Specifically, KGBugLocator first exploited interrelation information via code knowledge graph. Subsequently, it employed a keywords supervised bi-directional attention mechanism to mine the interaction information between bug reports and source code files. Ultimately, a fault localization component calculated similarity scores based on the integrated representations.
@@
Liang et al. [liang2022modeling] proposed FLIM, which leveraged CodeBERT to capture function-level code semantics and computed bug report-source code similarity through aggregated function interactions. FLIM consists of two stages. In the first stage, they initially built a training dataset called MIX dataset by splitting a source code file into a set of functions. Following this, they utilized the MIX dataset to fine-tune the CodeBERT model to make it adaptive to fault localization task. In the second stage, they employed the fine-tuned CodeBERT model to predict the relevance between the functions of the source code files and two parts (i.e., summary and description) of the bug report. Finally, semantic features extracted from CodeBERT and IR features are fused to a learning-to-rank model which gives the final relevance score.
@@
Han et al. [han2023bjxnet] proposed a multi-modal representation learning framework bjXnet, which integrated the text semantic features of the source code (shallow semantic features) and the graph features of the CPG (deep semantic features) by using the attention mechanism. Specifically, bjXnet employed the GatedGraphConv layer based on the GGNN to build the GatedGCN network for graph feature extraction, and utilized TextCNN as the text feature extractor to extract the text content of bug reports and source code files. Finally, bjXnet embedded bug reports, source code text, and source code CPGs into the same semantic space, comparing the relationship between report representation vectors and code representation vectors by the cosine similarity to determine whether the report and source code are related.
@@
Apart from focusing solely on semantic matching techniques, some researchers have embarked on endeavors to improve the performance of fault localization by exploring other perspectives and methods combined with advanced semantic models. For instances, Zhu et al. [zhu2022bl] proposed BL-GAN, which leveraged the adversarial training strategy and reinforcement learning for fault localization. In specific, the generator in their model generated code file paths for each given bug report, so that it can further construct pairs of bug report and code file. On the other hand, the discriminator attempted to distinguish whether the pair of bug report and code file is from the generator or not. BL-GAN adopted an attention-based Transformer architecture to capture semantic and sequence information to process bug reports, while incorporated a novel multi-layer GCN to process the source code in a graphical view to capture the proprietary structural information in code files, thereby enhancing the performance of fault localization.